What is the molecular mass (Formula mass) if a 3.2 g of the gas occupies a volume of 1.6 L at 30 degree Celsius and 745 mm Hg?


1
Expert's answer
2021-01-29T08:24:18-0500

PV = nRT

0,98 × 1,6 = n× 8,31 × 303,15

1,568 = n × 2519

n = 0,0006mol

M= 3,2 / 0,0006 = 5333g/mol